# Holographic Extended Thermodynamics of deformed AdS-Schwarzschild black hole

### v2

@@ 2.3 Critical Exponents
To compute the universal exponent $\alpha$, we utilize the entropy relation given in Eq. (12). Since the entropy is independent of temperature near the critical point, it follows that $\alpha=0$. The computation of the universal critical exponents for the RN-AdS black hole, corresponding to the case with control parameter $\beta=0$, has already been carried out in ref. Kubiznak:2012wp . However, the case with both parameters $\alpha$ and $\beta$ nonzero is non-trivial and necessitates a numerical treatment. We know for vdW phase, the equation of state is given by
@@
we plot the $\lnln\ln$ curve of $\kappa_{T}P_{c}$ versus (1T^)(1-$\hat${T}) along the $\hat{P}$$\hat{v}$ coexistence line, as shown in Fig. 10. We get $\gamma^{\prime}\approx 1$. Furthermore, from the scaling relation $\gamma=\beta(\delta-1)$, we obtain $\delta=3$.
@@
3 Extended boundary thermodynamics
@@ 3.1 Thermodynamic quantities for dual CFT
    $E=\frac{2Cx}{R}\left(1+x^{2}+\frac{\tilde{\alpha}(b^{2}+3x^{2}+3bx)}{3x(b+x)^{3}}\right),$ (44)
@@ 3.2 Thermodynamic ensembles in the dual CFT
$\text{Canonical ensemble, fixed }(\mathcal{V},C):\quad$ $F\equiv E-TS,$
$\text{Fixed }(p,C):\quad$ $G\equiv E-TS+p\mathcal{V},$
$\text{Fixed }(p,\mu):\quad$ $\Phi\equiv E-TS+p\mathcal{V}-\mu C,$
$\text{Fixed }(\mathcal{V},\mu):\quad$ $\Omega\equiv E-TS-\mu C.$
@@ 3.2.1 Phase transition in $(\mathcal{V},C)$
In this ensemble, the thermodynamic variables $\mathcal{V}$ and $C$ are held fixed, thereby defining the ensemble characterized by the pair $(\mathcal{V},C)$. We now proceed to investigate the behaviour of the free energy $F$ as a function of temperature $T$, for various fixed values of $(\mathcal{V},C)$. The corresponding free energy is defined as F=ETSF=E-TS and its expression is given by
$F=E-TS=\frac{2Cx}{R}\left(1+x^{2}+\frac{\tilde{\alpha}(b^{2}+3x^{2}+3bx)}{3x(b+x)^{3}}\right)-\left(\frac{1+x^{2}\left(3-\frac{\tilde{\alpha}}{(b+x)^{4}}\right)}{4\pi xR}\right)(4\pi Cx^{d-2}),$ (47)
@@
Finally, in the special case where the deformation parameter vanishes, i.e., $\tilde{\alpha}=0$, we recover a Hawking–Page-like phase transition. This behaviour is depicted in Fig 15. Although this transition was originally identified by Hawking Hawking:1982dh , it has since been recognized that the generalized Hawking–Page transition also occurs in various other black hole spacetimes Cong:2021jgb ; Tarrio:2011de ; Pedraza:2018eey . Furthermore, in the limit of a vanishing control parameter—corresponding to the CFT dual of the RN–AdS black hole—we observe a van der Waals–like phase structure, consistent with the results reported in Cong:2021jgb .
@@ 3.2.2 Phase transition in ensemble $(p,C)$
$G=E-TS+p\mathcal{V}=(TS+\mu C)-TS+p\mathcal{V}=\mu C+p\mathcal{V}\,.$ (52)
@@
The Gibbs free energy–temperature plot exhibits subtle and potentially novel features, with its structure varying significantly depending on the choice of parameters. In Fig 17 (Left), we have drawn three distinct GTG-T plots in which the green curve qualitatively differs from the curves in Fig 16. For a suitable set of parameters, the plot displays swallowtail-like behaviour characterised by three distinct branches—low-entropy (LE), intermediate-entropy (IE), and high-entropy (HE).
@@ 3.2.3 Phase transition in ensemble $(p,\mu)$
For the ensemble with fixed $(p,\mu)$, the corresponding thermodynamic potential is given by $\Phi=E-TS+p\mathcal{V}-\mu C$. Invoking the Euler relation, we find that the free energy simplifies to $\Phi=p\mathcal{V}$. Hence, in this ensemble, the thermodynamic potential is determined solely by the pressure and volume.
@@ 4 Conclusions and Discussions
In this work, we have investigated the thermodynamic phase transition of the deformed AdS-Schwarzschild black hole which is obtained from the gravitational decoupling (GD) method Khosravipoor:2023jsl , and followed by the study of phase transition of the thermal CFT states dual to the deformed AdS-Schwarzschild black hole in an extended phase space- treating the cosmological constant and its conjugate quantity as thermodynamic variables associated with the pressure and volume, respectively.
On the bulk side, the thermodynamic behaviour of the deformed AdS-Schwarzschild black hole is critically regulated by the deformation parameter $\alpha$ in a canonical ensemble, with distinct phase transitions emerging under specific conditions. When the control parameter $\beta\neq 0$, numerical analysis reveals two critical values, $\alpha_{c_{1}}$ and $\alpha_{c_{2}}$, bounding an interval $\alpha\in(\alpha_{c_{1}},\alpha_{c_{2}})$ where the system exhibits a first-order van der Waals (vdW) phase transition. This is characterized by a non-monotonic temperature profile (Fig. 1, left panel) and the canonical swallowtail structure in the PPvv diagram (Fig. 2). Outside this interval ($\alpha<\alpha_{c_{1}}$ or $\alpha>\alpha_{c_{2}}$), no vdW transition is observed. Moreover, when the deformation is turned off, we observe no vdW phase structure (Fig. 3, right panel). In the limit $\beta=0$, the solution reduces to a charged AdS black hole, where $\alpha$ acts as the square of the electric charge ($\alpha\propto Q^{2}$). Here, a single critical value $\alpha_{c}=6.25$ triggers a standard vdW transition (Fig. 3, left panel), consistent with the Reissner-Nordström-AdS universality class. Conversely, when $\alpha<\alpha_{c_{1}}$ (with $\beta\neq 0$), the system undergoes a Hawking-Page-type transition between thermal radiation and black hole phases (Fig. 6). Near criticality, the critical exponents universally satisfy $\alpha=0$, $\beta=1/2$, $\gamma=1$, $\delta=3$, matching mean-field theory predictions for vdW system.
In the dual boundary side, we first obtain the various thermodynamical quantities on the dual CFT states by assuming the variation of the cosmological constant. At the same time we allow the central charge $C$ to vary. The boundary thermodynamics of the dual conformal field theory exhibits distinctive phase behaviour governed by the dimensionless deformation parameter $\tilde{\alpha}=\alpha/l^{2}$. In the fixed-volume and central-charge ensemble $(\mathcal{V},C)$, a Hawking-Page transition occurs when $\tilde{\alpha}<\tilde{\alpha}_{c}$ for fixed $b=\beta/l$, manifesting as characteristic feature in free energy $F$ (Fig. 12). Increasing $\tilde{\alpha}$ elevates both $F$ and the Hawking temperature $T_{\rm HP}$ (Fig. 12, right panel), while increasing $b$ suppresses these quantities (Fig. 12, left panel). In the undeformed limit ($\tilde{\alpha}\to 0$), we recover the standard AdS-Schwarzschild result $T_{\rm HP}=(\pi R)^{-1}$. Notably, other ensembles show no conventional phase transitions: The ensemble $(p,C)$ exhibits no van der Waals-like behavior in Gibbs free energy $G$ (Figs. 1617), with stable phases confined exclusively to intermediate-entropy branches, while the ensemble $(p,\mu)$ does not show any interesting thermodynamic behavior in thermodynamic potential $\Phi=p\mathcal{V}$ (Fig. 18).
@@
From the bulk perspective, the deformation parameter $\alpha$ directly modulates gravitational geometry, creating a rich phase structure. This bulk behaviour holographically maps to confinement/deconfinement transitions in the boundary $(\mathcal{V},C)$ ensemble, where $\tilde{\alpha}<\tilde{\alpha}_{c}$ triggers deconfinement (Fig. 12 or 13). Here the GD deformation introduces new scales ($\alpha,\beta$) that independently control confinement regimes. Crucially, the deformation induces a mass gap in the dual theory spectrum, breaking conformal invariance and generating confinement dynamically through the $\beta$-regulated singularity avoidance at $r=0$. Such GD couplings represent controlled departures from Einstein gravity that preserve (weak) energy conditions while enriching phase diagrams. Future work should explore higher-dimensional generalizations, particularly since critical exponents exhibit $d$-dependent scaling in analogous deformations Gheorghiu:2013jha ; Estrada:2018vrl , suggesting dimensionally tunable phase structures. An interesting direction for future work involves examining other versions of gravitational decoupling approaches Maurya:2022wwa ; Brown:2017wpl and their potential impact on extended holographic thermodynamics.
@@ Acknowledgment
B.S would like to acknowledge Mr. Sukhdeep Singh Gill for his valuable discussions.

### v3

@@ 3.3 Critical Exponents
To compute the universal exponent $\alpha$, we utilize the entropy relation given in Eq. (13). Since the entropy is independent of temperature near the critical point, it follows that $\alpha=0$. The computation of the universal critical exponents for the RN-AdS-like black hole, corresponding to the case with control parameter $\beta=0$, has already been carried out in ref. Kubiznak:2012wp . However, the case with both parameters $\alpha$ and $\beta$ nonzero is non-trivial and necessitates a numerical treatment. We know for vdW phase, the equation of state is given by
@@
we plot the $\lnln\ln$ curve of $\kappa_{T}P_{c}$ versus (1T^)(1-$\hat${T}) along the $\hat{P}$$\hat{v}$ coexistence line, as shown in Fig. 10. We get $\gamma^{\prime}\approx 1$. Furthermore, from the scaling relation $\gamma=\beta(\delta-1)$, we obtain $\delta\approx 3$.
@@
In the following section, we analyse the boundary thermodynamics. We first construct the dual CFT metric and then examine the associated thermodynamic properties.
4 CFT thermodynamics with a chemical potential for the central charge
@@ 4.1 Thermodynamic quantities for dual CFT
    $\mathcal{E}=\frac{2Cx}{R}\left(1+x^{2}+\frac{\tilde{\alpha}(b^{2}+3x^{2}+3bx)}{3x(b+x)^{3}}\right),$ (45)
@@ 4.2 Thermodynamic ensembles in the dual CFT
$\text{Canonical ensemble, fixed }(\mathcal{V},C):\quad$ $F\equiv\mathcal{E}-TS,$
$\text{Fixed }(p,C):\quad$ $G\equiv\mathcal{E}-TS+p\mathcal{V},$
$\text{Fixed }(p,\mu):\quad$ $\Phi\equiv\mathcal{E}-TS+p\mathcal{V}-\mu C,$
$\text{Fixed }(\mathcal{V},\mu):\quad$ $\Omega\equiv\mathcal{E}-TS-\mu C.$
@@ 4.2.1 Phase transition in $(\mathcal{V},C)$
In this ensemble, the thermodynamic variables $\mathcal{V}$ and $C$ are held fixed, thereby defining the ensemble characterized by the pair $(\mathcal{V},C)$. We now proceed to investigate the behaviour of the free energy $F$ as a function of temperature $T$, for various fixed values of $(\mathcal{V},C)$. The corresponding free energy is defined as $F=\mathcal{E}-TS$ and its expression is given by
$F=\mathcal{E}-TS=\frac{2Cx}{R}\left(1+x^{2}+\frac{\tilde{\alpha}(b^{2}+3x^{2}+3bx)}{3x(b+x)^{3}}\right)-\left(\frac{1+x^{2}\left(3-\frac{\tilde{\alpha}}{(b+x)^{4}}\right)}{4\pi xR}\right)(4\pi Cx^{d-2}),$ (48)
@@
Finally, in the special case where the deformation parameter vanishes, i.e., $\tilde{\alpha}=0$, we recover a Hawking–Page-like phase transition. This behaviour is depicted in Fig 15. Although this transition was originally identified by Hawking Hawking:1982dh , it has since been recognized that the generalized Hawking–Page transition also occurs in various other black hole spacetimes Cong:2021jgb ; Tarrio:2011de ; Pedraza:2018eey . Furthermore, in the limit of a vanishing control parameter—corresponding to the CFT dual of the charge-like AdS black hole—we observe a van der Waals–like phase structure, consistent with the results reported in Cong:2021jgb .
@@ 4.2.2 Phase transition in ensemble $(p,C)$
$G=\mathcal{E}-TS+p\mathcal{V}=(TS+\mu C)-TS+p\mathcal{V}=\mu C+p\mathcal{V}\,.$ (53)
@@
The Gibbs free energy–temperature plot exhibits subtle and potentially novel features, with its structure varying significantly depending on the choice of parameters. In Fig 17 (Left), we have drawn three distinct GTG-T plots in which the green curve qualitatively differs from the curves in Fig 16. For a suitable set of parameters, the plot displays three distinct branches—low-entropy (LE), intermediate-entropy (IE), and high-entropy (HE). However, in the next section, the specific heat capacities plots reveal that they do not match the traditional plot of the swallowtail-like curve.
@@ 4.2.3 Phase transition in ensemble $(p,\mu)$
For the ensemble with fixed $(p,\mu)$, the corresponding thermodynamic potential is given by $\Phi=\mathcal{E}-TS+p\mathcal{V}-\mu C$. Invoking the Euler relation, we find that the free energy simplifies to $\Phi=p\mathcal{V}$. Hence, in this ensemble, the thermodynamic potential is determined solely by the pressure and volume.
@@
In the next section, we compute the critical points and examine the thermodynamic stability.
@@ 5 Conclusions
In this work, we have done an extensive investigation of the thermodynamic phase transition of the deformed AdS-Schwarzschild black hole which is obtained from the gravitational decoupling (GD) method Khosravipoor:2023jsl , and followed by the study of phase transition of the thermal CFT states dual to the deformed AdS-Schwarzschild black hole in an extended phase space, treating the cosmological constant and its conjugate quantity as thermodynamic variables associated with the pressure and volume, respectively.
In the analysis of bulk thermodynamics, we have encountered a less trivial two-critical-point structure in the temperature profile, corresponding to two distinct deformation-controlled critical values $\alpha_{c_{1}}$ and $\alpha_{c_{2}}$. Furthermore, the range $\alpha_{c_{1}}<\alpha<\alpha_{c_{2}}$ supports van der Waals–like transitions absent in both the Schwarzschild-AdS and RN–AdS limits. This thermodynamic behaviour is further supported by computing the universal ratio at the critical point; which is found to be close to the van der Waals value, and by confirming that the mean-field critical exponents are satisfied. On the dual CFT side, we find that the thermodynamic behaviour exhibits a Hawking–Page–like transition in the fixed $(\mathcal{V},\mathcal{C})$ ensemble, reflecting the confinement–deconfinement phase transition characteristic of the dual CFT. Additional ensembles were also considered; notably, the fixed $(p,\mathcal{C})$ ensemble displays unconventional thermodynamic phase structure, warranting further theoretical investigation.
@@
From the bulk perspective, the deformation parameter $\alpha$ directly modulates gravitational geometry, creating a rich phase structure. This bulk behaviour holographically maps to confinement/deconfinement transitions in the boundary $(\mathcal{V},C)$ ensemble, where $\tilde{\alpha}<\tilde{\alpha}_{c}$ triggers deconfinement. Here the GD deformation introduces new scales ($\alpha,\beta$) that independently control confinement regimes. Such GD couplings represent controlled departures from Einstein gravity that preserve (weak) energy conditions while enriching phase diagrams. Although our analysis focuses on the minimal geometric deformed background, a particular realization of the GD method, similar thermodynamic behavior may also arise in other GD frameworks. An interesting direction for future research would be to explore alternative formulations of gravitational decoupling Maurya:2022wwa ; Brown:2017wpl and their potential impact on extended holographic thermodynamics.
Future work should explore higher-dimensional generalizations, particularly since critical exponents exhibit $d$-dependent scaling in analogous deformations Gheorghiu:2013jha ; Estrada:2018vrl , suggesting dimensionally tunable phase structures.
@@ Acknowledgment
BS acknowledges support from the IIT Kharagpur Institute Fellowship. BS would like to also acknowledge Mr. Sukhdeep Singh Gill for his valuable discussions.
Appendix A Appendix: Smarr relation and deformation conjugates
In this appendix we compute the deformation conjugates $\Psi_{\alpha},\Psi_{\beta}$ for the mass function used in the main text and verify the Smarr relation explicitly. The mass as a function of the horizon radius $r_{h}$ is
$M(r_{h})=\frac{1}{6}\Big(3r_{h}+\frac{3r_{h}^{3}}{L^{2}}+\alpha\frac{\beta^{2}+3\beta r_{h}+3r_{h}^{2}}{(\beta+r_{h})^{3}}\Big).$ (62)
Deformation conjugates
The deformation conjugates $\Psi_{\alpha},\Psi_{\beta}$ are defined as partial derivatives of $M$ with respect to the parameters,
$\Psi_{\alpha}\;:=\;\Big(\frac{\partial M}{\partial\alpha}\Big)_{r_{h},L,\beta},\qquad\Psi_{\beta}\;:=\;\Big(\frac{\partial M}{\partial\beta}\Big)_{r_{h},L,\alpha}.$ (63)
Differentiating (62) gives the closed forms
$\Psi_{\alpha}$ $=\frac{1}{6}\,\frac{\beta^{2}+3\beta r_{h}+3r_{h}^{2}}{(\beta+r_{h})^{3}},$ (64)
$\Psi_{\beta}$ $=\frac{\alpha}{6}\,\frac{d}{d\beta}\!\left[\frac{\beta^{2}+3\beta r_{h}+3r_{h}^{2}}{(\beta+r_{h})^{3}}\right]=-\frac{\alpha\big(\beta^{2}+4\beta r_{h}+6r_{h}^{2}\big)}{6(\beta+r_{h})^{4}}.$ (65)
Thermodynamic quantities
The horizon equation $F(r_{h})=0$ is satisfied by the mass (62). The Hawking temperature, computed from the surface gravity, is
$T_{H}\;=\;\frac{F^{\prime}(r_{h})}{4\pi}\;=\;\frac{1}{4\pi r_{h}}\Bigg[1+r_{h}^{2}\!\Big(\frac{3}{L^{2}}-\frac{\alpha}{(\beta+r_{h})^{4}}\Big)\Bigg].$ (66)
We use the geometric entropy and thermodynamic volume
$S\;=\;\pi r_{h}^{2},\qquad V\;=\;\frac{4\pi}{3}r_{h}^{3},$ (67)
and the pressure defined by the cosmological constant (AdS4 bulk)
$P\;=\;\frac{3}{8\pi L^{2}}.$ (68)
Smarr relation (Euler homogeneity)
Under a global length rescaling $r_{h}\to\lambda r_{h}$, $L\to\lambda L$, $\beta\to\lambda\beta$ and $\alpha\to\lambda^{2}\alpha$ the mass scales as $M\to\lambda M$ (mass has length dimension one in four spacetime dimensions). Euler’s theorem for homogeneous functions therefore yields
$\begin{split}M&=2\;S\!\left(\frac{\partial M}{\partial S}\right)-2\;P\!\left(\frac{\partial M}{\partial P}\right)+2\;\alpha\!\left(\frac{\partial M}{\partial\alpha}\right)+1\;\beta\!\left(\frac{\partial M}{\partial\beta}\right)\\ &=2T_{H}S-2PV+2\alpha\Psi_{\alpha}+\beta\Psi_{\beta},\end{split}$ (69)
where we used the thermodynamic identifications $(\partial M/\partial S)_{P,\alpha,\beta}=T_{H}$ and $(\partial M/\partial P)_{S,\alpha,\beta}=V$.
Explicit Smarr check
Substituting (62)–(68) and (64)–(65) into the right-hand side of (69) and simplifying algebraically reproduces the left-hand side precisely. Thus the Smarr identity is satisfied exactly for the mass (62):
$M(r_{h})\;=\;2T_{H}S-2PV+2\alpha\Psi_{\alpha}+\beta\Psi_{\beta}.$ (70)
(We verified the equality symbolically on Mathematica; the algebraic cancellations are straightforward but somewhat long, and we therefore omit the intermediate expanded polynomial.)
Declaration
During the preparation of this work the authors used ChatGPT in order to provide me the useful information. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.